OpenStudy (anonymous):

Solve the equation 4 over 5 y + 2 = 6 for y. Please show your work and solution in the space provided

OpenStudy (binary3i):

4y/5+2=6 .............is this the equation

OpenStudy (anonymous):

please use bracket so we can understand a little bit ;)

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }y + 2 = 6\]

OpenStudy (binary3i):

y=5

OpenStudy (anonymous):

Will you explain?

OpenStudy (anonymous):